Two years ago a father was four times as old as his son. Three years from now the father will only be three times as old as his son. How old is each now?
** I've tried:
TWO YEARS AGO
Father = 4x - 2
Son = x - 2
I know the answer is {12, 42} But I still need to know how to execute the problem. I cant seem to come up with the right equation for "3 years from now", everything I tried ends up a negative number or just wrong.

Answer by ankor@dixie-net.com(12695)

Write an equation for each statement:
:
"Two years ago a father was four times as old as his son."
f - 2 = 4(s - 2)
simplify
f - 2 = 4s - 8
f = 4s - 8 + 2
f = 4s - 6
:
" Three years from now the father will only be three times as old as his son."
f + 3 = 3(s + 3)
f + 3 = 3s + 9
f = 3s + 9 - 3
f = 3s + 6
:
How old is each now?
From the 1st simplified equation, replace f in the above equation with (4s-6)
4s - 6 = 3s + 6
4s - 3s = 6 + 6
s = 12
:
I'm sure you get the idea now, you were on the right track.